OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. 2x - y = 7 2y = 4x - 14

Directrix (directrix):

2x - y = 7 2y = 4x - 14 Line up like variables.

OpenStudy (anonymous):

Would it be 1 solution?

Directrix (directrix):

2x - y = 7 -4x +2y = -14 --------------

Directrix (directrix):

@kbriones0 Multiply the top equation by 2 2x - y = 7 --> by 2 is what?

OpenStudy (anonymous):

14?

Directrix (directrix):

No. 2x - y = 7 --> by 2 ---> 4x -2y = -14 Add that to the second equation

OpenStudy (anonymous):

ok im with you

Directrix (directrix):

4x -2y = 14 -4x +2y = -14 ------------- Add those two equations. What do you get.

Directrix (directrix):

>Would it be 1 solution? NO

Directrix (directrix):

Look at the attached chart after you add the two equations.

OpenStudy (anonymous):

8x-4y-14=0?

Directrix (directrix):

No. 0 + 0 = 0 4x -2y = 14 is the same line as -4x +2y = -14. If you multiply -4x +2y = -14 by -1, you will get the first equation. So, there is only one line here. Or, you can say that they coincide in every point.

Directrix (directrix):

Your equations intersect in every point. Look on the attached chart and read how many solutions there are. That is the answer to this question.

OpenStudy (anonymous):

Oh okay. Thank you so much

Directrix (directrix):

You are welcome.